# Designing Adaptive User Interfaces for mHealth applications targeting chronic disease: A User-Centric Approach

### v1

@@ 5.1. Survey participants
| * Odds Ratios/(P-value),(CC=Content complexity, AD=Add on functions, DP=Different persuasive strategy, MI=Multimodal interaction) | |||||
5.2. Different types of adaptations
In our earlier SLR on AUIs for chronic disease support (Wang et al., 2023b), we compiled a comprehensive list of adaptations implemented in chronic disease-related applications by researchers. Our aim with this survey is to discern the users’ perspectives regarding the significance and value attributed to these adaptations. Our survey revealed a diversity of preferences among the respondents, with no singular adaptation type dominating over others. Content complexity (59%) and graphic design (58%) emerged as the most prevalent, closely followed by interface elements rearrangement (47%) and multimodal interaction (46%). In contrast, the sound effects exhibited the lowest frequency among the listed adaptations (28%) (see Table 5). This finding is not consistent with prevailing trends in existing AUI studies within the domain of chronic diseases (Wang et al., 2023b). Our SLR underscored graphic design (presentation adaptation) as the predominant adaptation type, which encompassed layout, font size, colour, and theme adjustments, while other adaptation types received comparatively less attention. However, our survey findings highlighted users’ significant emphasis on content adaptation and some types of behaviour adaptation. The difference in research trends and user preferences offers valuable information on the types of adaptation preferred by users. It underscores the importance of aligning researchers’ efforts with user preferences and priorities to ensure the effective design and implementation of adaptive user interfaces in chronic disease management applications.
@@ 5.2.1. Relationship between users’ preferences for adaptations and their demographic characteristics
We conducted a Chi-square independence test to investigate the influence of demographic variables on user preferences for adaptation types. The results indicate significant associations between age, nationality, chronic diseases, and preferences for adaptations such as content complexity, add-on functions, persuasive strategy, and multimodal interaction. Subsequently, a binary logistic regression analysis was employed to examine the relationships between various adaptations and significant demographic factors identified in the Chi-square test (see Table 6). It is important to note that statistical models such as binary logistic regression are built on certain assumptions about the data, including independent observations and no perfect multicollinearity (Harris, 2021), which are met in our dataset. Compared to participants from Australia, individuals from China are less likely to prefer adaptations related to content complexity (OR = 0.125) and add-on functions (OR = 0.215). Individuals from China may have different expectations or preferences regarding the complexity of content or additional functions in mHealth applications, leading to a lower likelihood of preferring content complexity adaptations compared to participants from Australia. Participants with cardiometabolic diseases showed a significantly higher inclination towards adaptations such as different persuasive strategies, additional functions, and content complexity compared to those with respiratory diseases. Similarly, participants with mental health conditions also demonstrated a notable preference for content complexity. These preferences may arise from the self-management nature of cardiometabolic and mental health conditions, which often necessitate individuals to actively monitor their health and follow treatment plans (Cruz-Ramos et al., 2022; Chow et al., 2016; Torous et al., 2015). Furthermore, participants over 45 years of age exhibited a higher tendency to seek multimodal interaction (OR = 5.824) and additional function (OR = 3.764) adaptations compared to their younger counterparts. This tendency among older users may stem from a preference to minimise interaction efforts and an increased need for additional assistance with app navigation (Trajkova and Martin-Hammond, 2020; Betts et al., 2019; Martin-Hammond et al., 2019). As individuals age, they may encounter challenges related to vision, dexterity or cognitive abilities, making features such as multimodal interaction and additional functions particularly appealing to facilitate their interaction with mHealth applications.
@@ 5.3. Data sources for adaptation
Our SLR identified various sources of data used as a basis for adaptation (Wang et al., 2023b). In our survey, we aimed to further explore users’ perspectives on the preferred types of data for adaptation. We found that participants exhibited various preferences about the source of adaptation data and almost half expressed the desire for the app to adapt based on physiological, physical, and psychological characteristics, user preferences, and user goals (see Table 5). The user characteristics data was the most popular compared to the interaction-related data and the task-specific data. This finding is consistent with existing research trends (Wang et al., 2023b). Interestingly, there was a notable preference for interaction-related and task-specific data in the prior research. This underscores the complexity of user preferences and the need for nuanced approaches to adaptation in user interfaces.
We used a Chi-square independence test to explore how demographic variables influence users’ preferences for data sources for adaptation. The findings revealed significant associations between nationality, chronic diseases, education level, and various types of data (see Table 7). Subsequently, we performed a binary logistic regression analysis to examine the relationships between different data sources and the significant results obtained from the Chi-square test. Participants from different countries exhibit varying preferences for data adaptation based on physiological, physical, preference, feedback, goals, and motivation factors. Given the relatively small number of participants from the UK and USA in our survey, it is essential to acknowledge the potential impact on the statistical power of our analysis. With a reduced sample size, there is a risk of diminishing the ability to detect genuine effects or associations accurately. Consequently, the observed relationships between nationality and data adaptation preferences may be less reliable, introducing uncertainty into our findings (Nemes et al., 2009). Individuals with higher levels of education exhibit a stronger inclination toward the adaptation of their motivation to use the application (Bachelor: OR= 7.835, Postgraduate: OR= 6.798). Individuals with higher levels of education often have a greater awareness of the benefits of using technology for health management and may also have a greater understanding of the importance of motivation in achieving health-related goals. Therefore, people with higher educational levels can exhibit a greater tendency to adapt their motivation to use the application due to their improved understanding of the role of motivation in achieving health outcomes (Woldaregay et al., 2018). Individuals with cardiometabolic diseases are more inclined to desire app adaptations that align with their goals (OR=13.948). This preference could be attributed to the management of cardiometabolic diseases, which often involves monitoring the levels of diet and physical activity and striving to achieve specific goals (Cruz-Ramos et al., 2022; Chow et al., 2016; Torous et al., 2015).
@@ 5.4. Preferred level of involvement in the adaptation
The SLR found varied ways of the involvement of the user in the adaptation process. In the survey, we discovered that participants generally favoured a semi-automatic adaptation approach (54%), which involves collaboration between the system and the end-users to achieve adaptation (Mukhiya et al., 2020; Abrahão et al., 2021). On the contrary, only a small minority of participants (8%) expressed a preference for a fully manual system, where users have complete control over the modification of specific UI elements to suit their needs (Akiki et al., 2014). The limited preference for fully manual systems suggests that users may find manual adaptation processes cumbersome or time-consuming, a sentiment that participants in the user study articulated and which was further corroborated by their feedback in the post-session survey. Fully manual systems may also require users to have a higher level of technical proficiency, which could be a barrier for some individuals, particularly those with limited technological literacy or cognitive abilities (Akiki et al., 2014). The SLR also indicated a lower preference for manual systems, which aligns with our survey findings. While significant research efforts have been dedicated to automatic systems, there has been a recent surge in interest in semi-automatic approaches (Wang et al., 2023b). No significant correlations are found between the level of involvement and demographic variables.
5.5. Key findings from survey
@@
Our survey offers valuable insights into users’ preferences regarding various aspects of design adaptations, which is partially consistent with existing research efforts. From our analysis, we have identified four key findings that help unravel the intricacies of user preferences. The survey results complement and provide context for the qualitative findings obtained from our interview and focus group studies.
@@
Key findings 1: Multimorbidity nature of chronic diseases. The prevalence of multimorbidity among participants underlines the common experience of individuals managing multiple chronic diseases simultaneously. This phenomenon is not unique to our study and has been consistently reported and emphasised in other research (Islam et al., 2014; Anderson et al., 2016). However, despite the widespread occurrence of multimorbidity, much of the existing exploration in the realm of mHealth applications for chronic disease management has predominantly focused on addressing a single chronic disease (e.g., hypertension (Hallberg et al., 2016) and asthma (Tinschert et al., 2017)). Similarly, in the context of AUIs, research has primarily concentrated on investigating their efficacy in managing single chronic diseases such as stroke (Burke et al., 2009) and diabetes (Pagiatakis et al., 2020). Consequently, there is a lack of approaches that address the diverse and complex requirements of individuals facing the challenges of managing multiple chronic conditions simultaneously. Closing this gap is essential to ensure that mHealth interventions and AUIs effectively support individuals in navigating the complexities associated with multimorbidity. This phenomenon links to what could be adapted in the UI and users’ preference for controllability. Users may feel overwhelmed by managing multiple diseases, leading to a lack of motivation and interest in actively intervening or taking responsibility themselves (see Figure 9).
@@
Key findings 2: Variations in purposes and usage pattern. The extensive use of mHealth applications for various purposes, ranging from symptom monitoring to dietary management, highlights their vital role in supporting the self-management of individuals with chronic diseases. Each app function serves a specific need, leading to different usage patterns among users. However, even within the same functionality, users may exhibit fluctuating usage patterns, with no fixed frequency or duration. Research indicates that individuals engage with physical activity tracking apps in intermittent intervals, characterised by periods of consistent use followed by breaks and subsequent reengagement (Lin et al., 2018; Meyer et al., 2016). These task-related characteristics have implications for the design of AUIs and their performance, especially concerning users’ preferences for controllability over adaptations and the timing and frequency of adaptation (Lavie and Meyer, 2010; Peissner and Edlin-White, 2013) (see Figure 9). For example, tasks that require minimal effort and are performed daily, such as recording mood, stress levels, or blood pressure, may not require manual control by the users. In such cases, users might prefer the system to handle these tasks automatically without requiring explicit confirmation. Additionally, since these tasks are performed regularly compared to other functionalities, users may not desire frequent changes to occur each time they log in. Detailed discussions on task characteristics and adaptation strategies can be found in Section 4.5.1.
@@
Key findings 3: Cultural and contextual differences. The analysis identified significant correlations between demographic variables such as age, nationality, and chronic diseases, and preferences for specific types of adaptations. For example, Chinese individuals exhibited a lower inclination toward content complexity adaptations compared to Australian participants, suggesting cultural and contextual differences in adaptation preferences. Moreover, individuals with higher levels of education demonstrated a stronger tendency to adapt their motivation to use the application, likely due to their deeper understanding of motivation’s role in health outcomes. Previous research highlights the importance of culturally specific design preferences in influencing system usefulness across different cultural contexts (Evers and Day, 1997). Various cultural factors significantly influence individuals’ preferences regarding UI perceptions and design. For example, religion plays a role in shaping users’ preferences for specific symbols and colours, influencing their implicit trust in the interface (Siala et al., 2004). Political orientation and social structure may affect users’ perception of the hierarchy and complexity of information presentation (Schmid-Isler, 2000). Individuals with different levels of education may require varying levels of support and cues from systems (Reinecke and Bernstein, 2013). When considering the actual adaptation process, cultural and contextual factors can significantly influence whether the adaptation meets the user’s expectations or not, potentially resulting in a mismatch with their cultural background. Additionally, user education levels can play a role in determining the ideal timing to initiate adaptation and the extent of support needed (see Figure 9).
@@
Key findings 4: Preference aligns with goal. Through our survey analysis, we discovered that different types of chronic disease can significantly influence the adaptations and data preferences of users. Specifically, participants with cardiometabolic diseases showed a greater inclination towards app adaptations that align with their usage goals. This preference likely arises from the pressing need for self-management inherent in cardiometabolic diseases. The availability of various applications for the management of chronic diseases or multimorbidities in app stores underscores the importance of adaptation aligning with users’ primary app usage goals and adding value to their overall app experience (Bricca et al., 2022). This resonates with the overarching categoryobserved in the user study section (see Figure 9), emphasising the essential value brought by adaptations and their alignment with users’ primary app usage goals (Trajkova and Martin-Hammond, 2020). Understanding these dynamics is crucial for developers and designers aiming to create adaptations for mHealth applications that not only support individuals with chronic diseases but are also valued by them. There are numerous options and elements available for adaptation on the UI, but it is essential to prioritise those that align with the user’s goals. However, selecting the perfect solution from such diverse options is challenging, as it depends on individual user experiences and the type of application and its adaptations, as suggested by the study conducted by Jameson and Schwarzkopf (2002). However, ongoing research continues to try to offer recommendations and guidance to users, helping them select the most appropriate options based on the data available within the system (Reinecke and Bernstein, 2013).
@@
6. Guidelines for AUI Design for mHealth applications targeting chronic disease
Our approach to developing our guidelines follows and adapts the framework proposed by Hermawati and Lawson (2015) and Quiñones et al. (2018). The methodology proposed by Quiñones et al. (2018) outlines various stages that facilitate the development process and allow the formal specification of guidelines through appropriate validations and refinements. However, some of these stages lack detailed explanations of their implementation. The involvement of users in the process and the expansion of guidelines within a particular domain are not sufficiently emphasised. On the other hand, Hermawati and Lawson (2015) have modified a user-centric approach to create usability heuristics tailored to specific domains. Consequently, we integrate elements from both methodologies into our guideline development process. The guideline developmrnt process is as follows (see Figure 10):
Exploratory stage: During our review of the literature, we gathered essential information to develop a new set of guidelines for designing AUIs in mHealth applications. This involved examining existing guidelines, design considerations, and recommendations pertinent to our research focus. An analysis of our literature review findings is presented in Section 6.1.
Experimental stage: We have conducted a user study comprising interviews, focus groups, and a survey to delve into users’ experiences, preferences, and challenges related to adaptive user interfaces in mHealth applications. Through the analysis and interpretation of these qualitative and quantitative data, we obtained valuable information to inform the creation of guidelines.
Descriptive stage: In the formalisation process of the main guidelines, we began by selecting and highlighting the most significant categories and concepts derived from all qualitative and quantitative data collected during the experimental stage. Initially, we grouped the needs identified from users and then aligned this information with the guidelines and design recommendations uncovered during the exploratory stage (Hermawati and Lawson, 2015). For any remaining information that cannot be directly mapped, we abstract new guidelines. Furthermore, we provide detailed information pertaining to the application domain, including definitions and illustrative examples to enhance clarity and applicability. The refined and finalised guidelines are available in Table 10.
Validation stage: Given the critical nature of this stage, which directly pertains to the specific domain, we prioritise the solicitation of feedback from the users we have in the interview and the focus group study, as recommended in the literature (Hermawati and Lawson, 2015). Further details regarding the validation process can be found in Section 6.2.

### v2

@@ 5.2. Different Types of Adaptations
| This table shows how demographic factors affect preferences for mHealth adaptations using binary logistic regression. • Each cell shows the odds ratio (OR) for preference likelihood and the p-value for statistical significance. • An OR indicates higher preference, OR indicates lower preference, and means statistical significance. | |||||
| * CC=Content complexity, AD=Add on functions, DP=Different persuasive strategy, MI=Multimodal interaction | |||||
@@ 5.2.1. Relationship between users’ preferences for adaptations and their demographic characteristics
Conducting a Chi-square independence test, we found significant associations between age, nationality, chronic diseases, and preferences for adaptations such as content complexity, add-on functions, persuasive strategy, and multimodal interaction. Subsequently, a binary logistic regression analysis was employed to examine the relationships between various adaptations and significant demographic factors identified in the Chi-square test (see Table 3). It is important to note that statistical models such as binary logistic regression are built on certain assumptions about the data, including independent observations and non-perfect multicollinearity (Harris, 2021), which are met in our dataset. Compared to participants from Australia, individuals from China are less likely to prefer adaptations related to content complexity (OR = 0.125) and add-on functions (OR = 0.215). Individuals from China may have different expectations or preferences regarding the complexity of content or additional functions in mHealth applications, leading to a lower likelihood of preferring content complexity adaptations compared to participants from Australia. Participants with cardiometabolic diseases showed a significantly higher inclination toward adaptations such as different persuasive strategies, additional functions, and content complexity compared to those with respiratory diseases. Similarly, participants with mental health conditions also demonstrated a notable preference for content complexity. These preferences may arise from the self-management nature of cardiometabolic and mental health conditions, which often necessitate individuals to actively monitor their health and follow treatment plans (Cruz-Ramos et al., 2022; Chow et al., 2016; Torous et al., 2015). Furthermore, participants over 45 years of age exhibited a higher tendency to seek multimodal interaction (OR = 5.824) and additional function (OR = 3.764) adaptations compared to their younger counterparts. This trend among older users may be due to a preference to minimize interaction efforts and an increased need for additional assistance with application navigation (Trajkova and Martin-Hammond, 2020; Betts et al., 2019; Martin-Hammond et al., 2019). As individuals age, they may face challenges related to vision, dexterity, or cognitive abilities, which makes features such as multimodal interaction and additional functions particularly appealing to facilitate their interaction with mHealth applications.
@@ 5.3. Data Sources for Adaptation
Our SLR identified various sources of data used as a basis for adaptation (Wang et al., 2023b). Based on these findings, the present survey aimed to investigate end-user preferences on the types of data they consider most appropriate to inform adaptation in mHealth applications. We found that the participants exhibited various preferences about the source of the adaptation data and almost half expressed the desire for the application to adapt according to physiological, physical, and psychological characteristics, user preferences, and user goals (see Table 2). The user characteristics data were the most popular compared to the interaction-related data and the task-specific data. This finding is consistent with existing research trends (Wang et al., 2023b).
@@
| This table shows how demographic factors affect preferences for using of different data source using binary logistic regression. • Each cell shows the odds ratio (OR) for preference likelihood and the p-value for statistical significance. • An OR indicates higher preference, OR indicates lower preference, and means statistical significance. | |||||||
@@
We used a Chi-square independence test to explore how demographic variables influence users’ preferences for data sources for adaptation. The findings revealed significant associations between nationality, chronic diseases, education level, and various types of data (see Table 4). Subsequently, we performed a binary logistic regression analysis to examine the relationships between different data sources and the significant results obtained from the Chi-square test. Participants from different countries exhibit varying preferences for data adaptation based on physiological, physical, preference, feedback, goals, and motivation factors. Given the relatively small number of participants from the UK and USA in our survey, it is essential to acknowledge the potential impact on the statistical power of our analysis. With a reduced sample size, there is a risk of diminishing the ability to detect genuine effects or associations accurately. Consequently, the observed relationships between nationality and data adaptation preferences may be less reliable, introducing uncertainty into our findings (Nemes et al., 2009). Individuals with higher levels of education exhibit a stronger inclination toward the adaptation of their motivation to use the application (Bachelor: OR= 7.835, Postgraduate: OR= 6.798). Individuals with higher levels of education often have a greater awareness of the benefits of using technology for health management and may also have a greater understanding of the importance of motivation in achieving health-related goals. Therefore, people with higher educational levels can show a greater tendency to adapt their motivation to use the application due to their improved understanding of the role of motivation in achieving health outcomes (Woldaregay et al., 2018). Individuals with cardiometabolic diseases are more inclined to desire adaptations that align with their goals (OR=13.948). This preference could be attributed to the treatment of cardiometabolic diseases, which often involves monitoring the levels of diet and physical activity and striving to achieve specific goals (Cruz-Ramos et al., 2022; Chow et al., 2016; Torous et al., 2015).
@@ 5.4. Preferred Level of Involvement in the Adaptation
The SLR found varied ways of the involvement of the user in the adaptation process (Wang et al., 2023b). In the survey, we discovered that participants generally preferred a mixed-initiative adaptation approach (54%), which involves collaboration between the system and the end-users to achieve adaptation (Mukhiya et al., 2020; Abrahão et al., 2021). In contrast, only a small minority of the participants (8%) expressed a preference for a fully manual system, where users have complete control over the modification of specific UI elements to suit their needs (Akiki et al., 2014). The limited preference for fully manual systems suggests that users may find manual adaptation processes cumbersome or time consuming, particularly those with limited technological literacy or cognitive abilities (Akiki et al., 2014). The SLR also indicated a lower preference for manual systems, which aligns with our survey findings. Although significant research efforts have been dedicated to automatic systems, there has been a recent surge in interest in mixed-initiative adaptation (Wang et al., 2023b). No significant correlations are found between the level of involvement and demographic factors.
5.5. Key Findings from the User Survey
@@
From our analysis, we have identified four key findings (KFs) that help unravel the intricacies of user preferences, complementing and contextualizing the qualitative insights from the interview and focus group studies (Figure 6).
KF1
@@
The multimorbidity nature of chronic diseases. The prevalence of multimorbidity among participants highlights the common experience of managing multiple chronic diseases simultaneously, a phenomenon consistently reported and highlighted in other research (Islam et al., 2014; Anderson et al., 2016). Despite the widespread occurrence of multimorbidity, much of the existing exploration in the realm of mHealth applications for chronic disease management has primarily focused on addressing a single chronic disease (e.g., hypertension (Hallberg et al., 2016) and asthma (Tinschert et al., 2017)). Similarly, in the context of AUIs, research has focused mainly on investigating their efficacy in managing single chronic diseases (Burke et al., 2009; Pagiatakis et al., 2020). This phenomenon links to what could be adapted in the UI and users’ preference for controllability. Users may feel overwhelmed by managing multiple diseases, leading to a lack of motivation and interest in actively intervening or taking responsibility themselves (see Figure 6).
Figure 6. Mapping between survey key findings and user study categories. -
KF2
@@
Variations in purpose and usage pattern. The extensive use of mHealth applications for various purposes, ranging from symptom monitoring to dietary management, highlights their vital role in supporting the self-management of individuals with chronic diseases. However, even within the same functionality, users may exhibit fluctuating usage patterns, with no fixed frequency or duration. Research indicates that individuals engage with physical activity tracking applications in intermittent intervals, characterized by periods of consistent use followed by breaks and subsequent reengagement (Lin et al., 2018; Meyer et al., 2016). These task-related characteristics have implications for the design of AUIs and their performance, especially concerning users’ preferences for controllability over adaptations and the timing and frequency of adaptation (Lavie and Meyer, 2010; Peissner and Edlin-White, 2013) (see Figure 6). For example, tasks that require minimal effort and are performed daily, users might prefer the system to handle these tasks automatically without requiring explicit confirmation and not desire frequent changes to occur each time they log in.
KF3
@@
Cultural and contextual differences. The analysis identified significant correlations between demographic factors such as age, nationality, and chronic diseases, and preferences for specific types of adaptation. For example, Chinese individuals exhibited a lower inclination toward content complexity adaptations compared to Australian participants, suggesting cultural and contextual differences in adaptation preferences. This observation could be explained by Hofstede’s cultural dimensions model (Hofstede, 2001), which is widely used to examine human-computer interaction and cross-cultural challenges in UI design. Specifically, Hofstede’s Uncertainty Avoidance dimension, which is defined as the degree to which individuals in a culture perceive ambiguity as a threat and seek to reduce it, offers an explanation. In societies with high uncertainty avoidance, individuals tend to favor predefined UI over experimenting with new adaptations, whereas a society with lower uncertainty avoidance may exhibit greater openness to try new adaptations (Alsswey and Al-Samarraie, 2021). Previous research highlights the importance of culturally specific design preferences in influencing system usefulness in different cultural contexts (Evers and Day, 1997). For example, political orientation and social structure may affect users’ perception of the hierarchy and complexity of information presentation (Schmid-Isler, 2000). Moreover, individuals with higher levels of education demonstrated a stronger tendency to adapt their motivation to use the application, likely due to their deeper understanding of motivation’s role in health outcomes. Individuals with different levels of education may require varying levels of support and cues from the systems (Reinecke and Bernstein, 2013) (see Figure 6). Cultural and contextual factors significantly influence how well an adaptation aligns with user expectations, and misalignment may result in cultural mismatches that hinder usability and acceptance.
KF4
@@
Preference aligns with goal. Participants with cardiometabolic diseases showed a greater inclination toward adaptations that align with their usage goals. This preference likely arises from the urgent need for self-management inherent in cardiometabolic diseases. The availability of various applications for the management of chronic diseases or multimorbidities in the App Store underscores the importance of adapting to user primary usage goals and adding value to their overall application experience (Bricca et al., 2022). This resonates with the overarching category observed in the user study section (see Figure 6), emphasizing the essential value brought by adaptation and its alignment with users’ primary usage goals (Trajkova and Martin-Hammond, 2020). Selecting the perfect solution from such diverse options is challenging, as it depends on individual user experiences and the type of application and its adaptations, as suggested by the study conducted by Jameson and Schwarzkopf (2002), while ongoing research continues to try to offer recommendations and guidance to users, helping them select the most appropriate options based on the data available within the system (Reinecke and Bernstein, 2013).
@@
6. Stage Two: Guidelines for designing AUIs in mHealth applications targeting chronic diseases
Our guideline development approach draws on and integrates the methodologies proposed by Hermawati and Lawson (2015) and Quiñones et al. (2018). While Quiñones et al. (2018) outlines structured stages for guideline specification and validation, their framework offers limited emphasis on implementation procedures and user involvement. In contrast, Hermawati and Lawson (2015) adopts a user-centric approach specifically tailored to develop domain-specific heuristics. By combining the strengths of both methodologies, we designed a hybrid process that incorporates iterative validation, domain relevance, and active end-user participation throughout the development of AUI guidelines.
Exploratory phase: Our review of the literature offered critical insights that informed the development of AUI design guidelines for mHealth applications. These guidelines are grounded in existing frameworks and design considerations identified in prior research. A detailed analysis supporting this development is presented in Section 6.1.
Experimental phase: User studies using interviews, focus groups, and a survey explored user experience, preference, and challenges with AUI in mHealth applications. The analysis of both qualitative and quantitative data from this study provided valuable insights that informed the development of the design guidelines.
Descriptive phase: To formalize the primary guidelines, key categories and concepts were derived from the qualitative and quantitative data gathered during the experimental phase. The user needs were systematically grouped and aligned with the guidelines and design recommendations established during the exploratory phase, following the approach proposed by Hermawati and Lawson (2015). For insights that did not correspond to existing recommendations, new guidelines were abstracted, accompanied by precise definitions of the application domain and illustrative examples to improve clarity, relevance, and practical applicability.
Validation phase: Given the critical importance of this phase, which is deeply embedded within the specific application domain, particular emphasis was placed on collecting feedback from end-users to ensure the relevance of the guidelines (Hermawati and Lawson, 2015). Building on this foundation, input was also obtained from software practitioners responsible for implementing these guidelines, allowing for further refinement based on their practical experience. Finally, to assess the applicability and effectiveness of the developed guidelines, a case study was conducted in which real-world mHealth applications were evaluated using the proposed guideline set.